OpenStudy (anonymous):

Can someone describe the relationship between the graph of f(x)=lnx and g(x)=e^x?

OpenStudy (anonymous):

Did you graph them?

OpenStudy (anonymous):

I don't want to use a calculator to do it, but I guess I might as well go ahead and do it. I just wanted to see if I can focus on key similarities and differences between the two without needing to resort to a calculator.

OpenStudy (anonymous):

Well, those there are inverses of each other.

OpenStudy (anonymous):

That is what they are looking for.

OpenStudy (anonymous):

y=e^x try solving for x. you get x=ln(x)

OpenStudy (anonymous):

Well, did you already know they were inverses?

OpenStudy (anonymous):

i mean x=ln(y)

OpenStudy (anonymous):

Okay. Oh! I feel like a total idiot. I didn't even notice that or think to even do that.

OpenStudy (anonymous):

Cause if you know that, then you know how they're related. But if you didn't know that then it's interesting to see how inverses are related.

OpenStudy (anonymous):

I'm completely stressing about this test and I'm not looking at the big picture, I'm just trying to make things harder than they are.

OpenStudy (anonymous):

:D It gets easier. When you are in calculus you will just get stuff like that in a snap. Of course, by that time you won't get homework about it... oh, the irony...

OpenStudy (anonymous):

Haha! Well, I'm at the end of my class, and I'm trying to brush up for the AP test...

OpenStudy (anonymous):

What class?

OpenStudy (anonymous):

alma, did you know e^x is the inverse of ln(x)?

OpenStudy (anonymous):

Not in the beginning, but he/she figured it out.

OpenStudy (anonymous):

Yes. I did and I just realized that I knew that. I was just trying to memorize properties which I don't need to because they are simply inverses.

OpenStudy (anonymous):

If you know that two functions are inverses, then they will be mirror images of eachother around the line y=x. That's how the two graphs are related.

OpenStudy (anonymous):

properties of the graph. Thank you all for your help.